# Uncertainty modeling for fine-tuned implicit functions

### v1

@@ 4.1 Dropsembles
In contrast, dropout involves training only a single model instance. Applying dropout to a neural network involves sampling a "thinned" network, effectively the same as applying a binary mask to the weights. However, unlike ensembles where each model is trained independently, dropout results in networks that are not independent; they share weights. Thus, training a neural network with dropout is akin to simultaneously training a collection of $2^{n}$ thinned networks, all sharing a substantial portion of their weights. However, often this comes at a price of less accurate predictions.
In Dropsembles, the model is first trained with dropout on dataset $\mathcal{D}_{A}$. Subsequently, $M$ thinned network $\{f_{\theta_{m}}\}_{m\in[M]}$ instances are generated by sampling binary masks. Each of these "thinned" instances is then fine-tuned on dataset $\mathcal{D}_{B}$ independently, effectively creating an ensemble of "thinned" networks initialized with correlated weights. For inference and uncertainty estimations, this ensemble is treated as a uniformly weighted mixture model, and the predictions are aggregated in the same manner as traditional ensembles: $p(y|x)=\frac{1}{M}\sum_{m=1}^{M}p_{\theta_{m}}(y|x,\theta_{m})$. Although ensembles typically benefit from networks being large and independent, our experiments show that this relaxation of independence does not significantly diminish prediction performance metrics or expected calibration error. The overall training procedure is summarized in Algorithm 1.
@@ 4.2 Elastic Weight Consolidation regularization for implicit shape modeling
In the basic version of optimization described above, networks start their training initialized from the learned posterior of dataset $\mathcal{D}_{A}$. However, there is no guarantee that during fine-tuning the network won’t diverge arbitrarily from initial weights. This is particularly problematic for implicit shape modeling, where it is essential to retain information from a large and dense prior dataset $\mathcal{D}_{A}$ while adapting to sparse and noisy data $\mathcal{D}_{B}$ to avoid overfitting. To address this concern, we borrow developments from the continual learning literature, as they seamlessly fit into the framework described above.
In particular, when fine-tuning individual instances of thinned networks on dataset $\mathcal{D}_{B}$, we can apply the same reasoning to each network instance as described in EWC. Thus, the learning objective for part B is replaced by the objective described in equation (2), which includes an additional regularization term.
@@ 5 Experimental results
Our objective is to provide trustworthy predictions on dataset B without hurting performance. We do so by modeling the uncertainty of the weights of a fine-tuned model. In standard prediction tasks, models are evaluated with respect to accuracy and Expected Calibration Error (ECE). However, our specific setup (reconstruction from sparse views) calls for additional evaluation metrics suitable to computer vision tasks, such as Dice Score Coefficient (DSC) and Hausdorff distance. Besides ECE, we include reliability diagrams [65] given our preference for a more conservative modeling approach. Details on evaluation metrics can be found in Appendix B.
@@ 5.1 Classification under distribution shift
We first demonstrate our method on a toy data set for binary classification. This allows quantitative analysis and demonstrates the versatility of our method. The experiment focuses on fine-tuning a trained model on few samples with distribution shift compared to the original training samples. We generated two-dimensional datasets A and B with a sinusoidal decision boundary and Gaussian noise. A moderate distribution shift was modeled between the datasets by adjusting the support values. We created 1000 train samples for dataset A, and 50 train samples for dataset B. Training details can be found in Appendix C.
All variations of Dropsembles and Ensembles were tested using just 4 network instances. The best regularization parameter was selected for each method via an ablation study in Appendix D. All metrics in Table 1 were averaged across 3 random seeds. From the evaluations (Table 1, Figure 2), it is apparent that EWC improves the performance of Ensembles and Dropsembles on both datasets.
Figure 2: Toy classification example. a) Training data for binary classification task ("red" vs "blue") from datasets A (points, light) and B (crosses, dark) b) MC dropout trained only on Dataset A. c) Comparison of methods fine-tuned on Dataset B. Points are colored by the predicted class.
Table 1: Comparison of methods fine-tuned on dataset B for toy classification and corrupted MNIST reconstruction experiments. The results are averaged over three samples. All metrics reported in [%].
Toy classification MNIST
Method Acc-A $\uparrow$ ECE-A $\downarrow$ Acc-B $\uparrow$ ECE-B $\downarrow$ DSC $\uparrow$ ECE $\downarrow$
MCdropout $59.0\pm 5.3$ $40.0\pm 4.3$ $90.5\pm 0.6$ $11.5\pm 1.7$ $58.2\pm 3.1$ $8.0\pm 1.1$
Dropsembles $59.2\pm 4.3$ $39.8\pm 4.3$ $90.5\pm 0.6$ $9.2\pm 1.3$ $64.9\pm 6.6$ $6.8\pm 1.4$
Ensembles $56.5\pm 2.4$ $42.5\pm 1.7$ $90.0\pm 0.0$ $8.5\pm 0.6$ $62.2\pm 3.5$ $6.5\pm 0.6$
MCdropout + EWC $66.8\pm 5.6$ $33.0\pm 4.8$ $86.5\pm 2.1$ $10.5\pm 1.9$ $60.5\pm 4.8$ $7.1\pm 1.5$
Dropsembles + EWC $\textbf{96.2}\pm 4.2$ $\textbf{5.8}\pm 5.3$ $\underline{93.5}\pm 3.8$ $\underline{7.0}\pm 1.4$ $\underline{70.3}\pm 3.9$ $\underline{6.4}\pm 2.5$
Ensembles + EWC $\underline{95.8}\pm 2.5$ $\underline{7.5}\pm 2.4$ $\textbf{95.5}\pm 1.3$ $\textbf{5.5}\pm 1.0$ $\textbf{71.3}\pm 3.6$ $\textbf{5.2}\pm 1.9$
5.2 Implicit shape modeling from sparse views
@@
In our next experiment, we explore the reconstruction from sparse inputs using the MNIST dataset. To this end, we converted the images into binary masks through thresholding. To simulate sparse input conditions, we applied a grid mask to the images, masking out every third row and column. This masking strategy was consistently applied across both datasets A and B. For dataset A, we utilized all images for a single digit "7" from the MNIST training set.
@@
To introduce a moderate distribution shift and mimic common real-world dataset corruptions, we rotated the images in dataset B by 15 degrees and obscured approximately 15 percent of the original pixels. In this section we are focusing on shape reconstruction, therefore "fine-tuning" and "testing" are applied per image and not per dataset. Given that the occupancy network is trained at the pixel level, each of these images effectively constitutes an individual "dataset B". We randomly selected three distinct images from the MNIST test split, which would give us three different variations of "dataset B". Fine-tuning and evaluating the whole test split of MNIST dataset would be computationally demanding, as each occupancy network is fine-tuned on an individual image.
@@
Our occupancy network comprises an 8-layer MLP, designed to process the latent representation along with the 2-dimensional coordinates of each pixel, thereby facilitating pixel-wise predictions. For configurations utilizing dropout, we incorporated a dropout layer with a probability of $p=0.3$ following each linear layer in the network. Training details can be found in Appendix C.
@@
We conducted an ablation study to identify the optimal regularization parameter for all methods, for which details are provided in the Appendix D. We demonstrate qualitative predictions and associated uncertainties for a single-image example in Figure 3. Visual inspection of the uncertainty estimates reveals that EWC-regularized ensembles exhibit the desired behavior: they not only deliver accurate predictions but also provide conservative uncertainty estimates, particularly noting high uncertainty in regions with data corruption. Reliability diagrams, along with quantitative evaluations presented in Table 1, confirm that elastic regularization enhances the performance across all methods. Notably, EWC-regularized Dropsembles achieve performance comparable to that of the Ensembles but with significantly reduced resource usage.
@@
Figure 3: Corrupted MNIST reconstruction example. a) Example of training images. b) Comparison of fine-tuned methods on dataset B. A "perfectly calibrated" method would have reliability diagrams aligned on the diagonal. A "good conservative" method would have all bars above the diagonal.
@@
In this section, we adapt the data preparation procedure from [11] but replace the high-quality CT dataset with a synthetic dataset of anatomical shapes. As an anatomical shape prior, we utilize a rigged anatomical model from "TurboSquid". To model patient-specific variability and variations in poses during MRI acquisition, we generated 94 rigged deformations. Point-cloud models from the atlas were converted to voxels at 256 voxel resolution in order to correspond to the resolution of the target MRI dataset. We further applied random elastic deformations to mimic patient-specific shape variability. We created a paired "sparse" - "dense" dataset by selecting a consistent set of 17–21 sagittal slices and additionally applied two iterations of connected erosions to simulate patient-specific automatic MR segmentations [11]. A bicubic upsampling was employed on the "sparse" inputs to adapt them for the encoder. We trained ReconNet [11] on the entire training split of the atlas dataset to obtain an encoder, which we subsequently kept frozen. For the implicit decoder, we employed only one of the rigged samples to accurately quantify performance on synthetic data and demonstrate the method’s robustness to medium and strong distribution shifts. For real-world applications, we recommend using the entire rigged transformations dataset.
For the occupancy network, we employed the MLP architecture described by [6], which features eight linear layers, each of 128 dimensions (described in Appendix C). The initial training of the occupancy network was conducted on an anatomical atlas. This intensive training process required 48 hours on a single NVIDIA RTX A6000 GPU, equipped with 48 GB of memory. Given the significant resource demands, conducting a full ensemble comparison was deemed impractical. Such an analysis would require the training of multiple network instances, substantially increasing computational resource consumption. It is important to highlight that all models were trained at a very high resolution, which is critical in medical applications where precision is paramount.
@@
As the target dataset B, we use a publicly available dataset from [66], which comprises MR+CT images from 20 subjects. We employed the same segmentation network as [11] to obtain segmentations of 5 vertebrae, 5 discs, and the spinal canal [67]. These automatically generated segmentations of high-quality MRI samples were used as the ground truth for the sparse 3D reconstruction task. To create sparse inputs, we removed the same set of sagittal slices as in the atlas dataset. We randomly selected 3 subjects for the consequent fine-tuning and testing.
The MR dataset described above lacks ground truth segmentations, which complicates numerical evaluation. The goal of the model is to impute missing or misclassified parts using anatomical atlas priors, which, while enhancing reconstruction, are typically marked as incorrect in standard segmentation metrics. To facilitate a more accurate assessment of reconstruction metrics, we introduce an intermediate benchmark. For this purpose, we utilize three random rigged deformations from the atlas that were not exposed during training and apply a doubled level of erosions. This strong augmentation provides a challenging input for dataset B and allows to model a significant distribution shift aimed at rigorously testing the model under adverse conditions.
Numerical evaluation in Table 2 demonstrates consistent improvement of our method upon MC-dropout. In order to investigate the performance of the proposed method on MR dataset, we perform a detailed qualitative analysis in Figure 4. A consistent and distinct pattern not captured by the metrics alone, stands out from the qualitative assessment: MC-dropout tends to produce noisy reconstructions, as illustrated in Figure 4c-d, and generally yields predictions that lack coherent, continuous shapes. This behaviour is partially captured in metric "DSC avg" in Table 2 - dice score evaluated on individual samples of the network instances. This observation is notable because, although models trained with conventional dropout generate reasonable predictions, their use in uncertainty estimation undermines the fundamental objective of 3D modeling. Samples drawn from the dropout distribution do not yield plausible shapes, as evidenced in Figure 4d. In contrast, our model, which draws from deep ensembles, does not exhibit these limitations. Furthermore, increasing the number of samples in MC-dropout does not resolve this issue but rather leads to higher computational demands, as documented in Table 2.
EWC did not show significant benefits in this experiment, despite its superior performance in previous examples. We believe the challenge lies in selecting the appropriate regularization strength, compounded by the heavy computational demands of high-resolution 3D reconstruction. Nevertheless, EWC did not degrade performance and can be applied safely even with suboptimal weights. This critical aspect of model tuning highlights a complex area that demands further investigation.
@@
Table 2: Comparison of methods fine-tuned on dataset B of lumbar spine experiment. Evaluations are performed on the corrupted atlas. Metrics are reported for dataset B. Baseline is trained on dataset A.
Method DSC [%] $\uparrow$ DSC avg [%] $\uparrow$ HD $\downarrow$ ECE [%] $\downarrow$ Inference [sec] $\downarrow$
Baseline $65.0\pm 2.0$ $63.9\pm 1.8$ $17.5\pm 1.2$ $25.2\pm 0.4$ 167
MCdropout $85.3\pm 2.4$ $83.0\pm 2.0$ $13.1\pm 2.6$ $\underline{3.4}\pm 1.8$ 164
MCdropout + EWC $85.3\pm 2.4$ $83.1\pm 2.0$ $12.3\pm 2.4$ $\textbf{3.3}\pm 1.8$ 117
Dropsembles $\underline{86.8}\pm 2.4$ $\underline{86.3}\pm 2.3$ $\underline{11.4}\pm 1.8$ $4.6\pm 1.7$ 22
Dropsembles + EWC $\textbf{86.9}\pm 2.4$ $\textbf{86.4}\pm 2.3$ $\textbf{10.9}\pm 2.1$ $4.5\pm 1.7$ 21

### v2

@@ 4.1 Dropsembles
In contrast, dropout involves training only a single model instance. Applying dropout to a neural network involves sampling a “thinned” network, effectively the same as applying a binary mask to the weights. However, unlike ensembles where each model is trained independently, dropout results in networks that are not independent; they share weights. Thus, training a neural network with dropout is akin to simultaneously training a collection of $2^{n}$ thinned networks ($n$ number of weights), all sharing a substantial portion of their weights. However, often this comes at a price of less accurate predictions.
In Dropsembles, the model is first trained with dropout on dataset $\mathcal{D}_{A}$. Subsequently, $M$ thinned network $\{f_{\theta_{m}}\}_{m\in[M]}$ instances are generated by sampling binary masks. Each of these “thinned” instances is then fine-tuned on dataset $\mathcal{D}_{B}$ independently, effectively creating an ensemble of “thinned” networks initialized with correlated weights. For inference and uncertainty estimations, this ensemble is treated as a uniformly weighted mixture model, and the predictions are aggregated in the same manner as traditional ensembles: $p(y|x)=\frac{1}{M}\sum_{m=1}^{M}p_{\theta_{m}}(y|x,\theta_{m})$. Although ensembles typically benefit from networks being large and independent, our experiments show that this relaxation of independence does not significantly diminish prediction performance metrics or expected calibration error. The overall training procedure is summarized in Algorithm 1.
@@ 4.2 Elastic Weight Consolidation regularization for implicit shape modeling
In the basic version of optimization described above, networks start their training initialized from the learned posterior of dataset $\mathcal{D}_{A}$. However, there is no guarantee that during fine-tuning the network will not diverge arbitrarily from initial weights. This is particularly problematic for implicit shape modeling, where it is essential to retain information from a large and dense prior dataset $\mathcal{D}_{A}$ while adapting to sparse and noisy data $\mathcal{D}_{B}$ to avoid overfitting. To address this concern, we borrow developments from the continual learning literature, as they seamlessly fit into the framework described above.
In particular, when fine-tuning individual instances of thinned networks on dataset $\mathcal{D}_{B}$, we can apply the same reasoning to each network instance as described in EWC. Thus, the learning objective for part B is replaced by the objective described in (Equation 2), which includes an additional regularization term.
@@ 5 Experimental results
Our objective is to provide trustworthy predictions on dataset B without hurting performance. We do so by modeling the uncertainty of the weights of a fine-tuned model. In standard prediction tasks, models are evaluated with respect to accuracy and Expected Calibration Error (ECE). However, our specific setup (reconstruction from sparse views) calls for additional evaluation metrics suitable to computer vision tasks, such as Dice Score Coefficient (DSC) and Hausdorff distance. Besides ECE, we include reliability diagrams (Guo et al., 2017b) given our preference for a more conservative modeling approach. Details on evaluation metrics can be found in Appendix B.
All models were evaluated using four network instances. Training details are provided in Appendix C. The optimal regularization parameter for EWC was determined through an ablation study presented in Appendix D. Additional ablation results on the number of network instances are included in Appendix E.
@@ 5.1 Classification under distribution shift
We first demonstrate our method on a toy data set for binary classification. We generated two-dimensional datasets A and B with a sinusoidal decision boundary and Gaussian noise. A moderate distribution shift was modeled between the datasets by adjusting the support values. We created 1000 train samples for dataset A, and 50 train samples for dataset B. All metrics in Table 1 were averaged across 3 random seeds. From the evaluations (Table 1, Figure 2), it is apparent that EWC improves the performance of Ensembles and Dropsembles on both datasets.
Figure 2: Toy classification example. a) Training data for binary classification task (“red” vs “blue”) from datasets A (points, light) and B (crosses, dark) b) MC dropout trained only on Dataset A. c) Comparison of methods fine-tuned on Dataset B. Points are colored by the predicted class. EWC consistently improves both accuracy and uncertainty estimates on both A and B datasets.
Table 1: Comparison of methods fine-tuned on dataset B for toy classification, MNIST and ShapeNet reconstruction experiments. Values are shown in percentages [%] for all metrics.
Toy classification MNIST ShapeNet
Method Acc-A $\uparrow$ ECE-A $\downarrow$ Acc-B $\uparrow$ ECE-B $\downarrow$ DSC $\uparrow$ ECE $\downarrow$ mDSC $\uparrow$ ECE $\downarrow$
MCdropout $59.0\pm 5.3$ $40.0\pm 4.3$ $90.5\pm 0.6$ $11.5\pm 1.7$ $54.8\pm 6.9$ $8.6\pm 2.0$ $82.8\pm 2.2$ $19.31\pm 3.97$
Dropsembles $59.2\pm 4.3$ $39.8\pm 4.3$ $90.5\pm 0.6$ $9.2\pm 1.3$ $64.1\pm 4.8$ $6.1\pm 2.7$ $83.5\pm 2.1$ $19.51\pm 4.64$
Ensembles $56.5\pm 2.4$ $42.5\pm 1.7$ $90.0\pm 0.0$ $8.5\pm 0.6$ $62.6\pm 5.1$ $\underline{5.8}\pm 2.3$ $82.6\pm 2.2$ $25.88\pm 3.27$
MCdropout + EWC $66.8\pm 5.6$ $33.0\pm 4.8$ $86.5\pm 2.1$ $10.5\pm 1.9$ $55.7\pm 7.2$ $8.4\pm 2.2$ $\underline{83.9}\pm 2.4$ $\textbf{16.96}\pm 3.81$
Dropsembles + EWC $\textbf{96.2}\pm 4.2$ $\textbf{5.8}\pm 5.3$ $\underline{93.5}\pm 3.8$ $\underline{7.0}\pm 1.4$ $\textbf{70.3}\pm 1.7$ $6.1\pm 1.9$ $\textbf{84.3}\pm 2.1$ $\underline{17.67}\pm 4.45$
Ensembles + EWC $\underline{95.8}\pm 2.5$ $\underline{7.5}\pm 2.4$ $\textbf{95.5}\pm 1.3$ $\textbf{5.5}\pm 1.0$ $\underline{69.3}\pm 3.8$ $\textbf{5.2}\pm 2.0$ $83.4\pm 1.9$ $24.59\pm 2.98$
5.2 Implicit shape modeling
@@
In our next experiment, we explore the reconstruction from sparse inputs using the MNIST dataset. To this end, we converted the images into binary masks through thresholding. To simulate sparse input conditions, we applied a grid mask to the images, masking out every third row and column. This masking strategy was consistently applied across both datasets A and B. For dataset A, we utilized all images for a single digit “7” from the MNIST training set.
@@
To introduce a moderate distribution shift and mimic common real-world dataset corruptions, we rotated the images in dataset B by 15 degrees and obscured approximately 15 percent of the original pixels. In this section we are focusing on shape reconstruction, therefore “fine-tuning” and “testing” are applied per image and not per dataset. Given that the occupancy network is trained at the pixel level, each of these images effectively constitutes an individual “dataset B”. We randomly selected 20 distinct images from the MNIST test split, which would give us 20 different variations of “dataset B”. Fine-tuning and evaluating the whole test split of MNIST dataset would be computationally demanding, as each occupancy network is fine-tuned on an individual image.
@@
Our occupancy network comprises an 8-layer MLP, designed to process the latent representation along with the 2-dimensional coordinates of each pixel, thereby facilitating pixel-wise predictions. For configurations utilizing dropout, we incorporated a dropout layer with a probability of $p=0.3$ following each linear layer in the network.
@@
We demonstrate qualitative predictions and associated uncertainties for a single-image example in Figure 3. Visual inspection of the uncertainty estimates reveals that EWC-regularized ensembles exhibit the desired behavior: they not only deliver accurate predictions but also provide conservative uncertainty estimates, particularly noting high uncertainty in regions with data corruption. Reliability diagrams, along with quantitative evaluations presented in Table 1, confirm that elastic regularization enhances the performance across all methods. Notably, EWC-regularized Dropsembles achieve performance comparable to that of the Ensembles but with significantly reduced resource usage.
@@
Figure 3: Corrupted MNIST reconstruction example. a) Example of training images. b) Comparison of fine-tuned methods on dataset B. A “perfectly calibrated” method would have reliability diagrams aligned on the diagonal. A “good conservative” method would have all bars above the diagonal.
ShapeNet
We demonstrate the applicability of Dropsembles for modeling uncertainty in SDFs on a commonly used ShapeNet dataset. For dataset A, we utilized 1,000 randomly selected airplane shapes. For dataset B, we selected 10 unseen airplane shapes and introduced a distribution shift by applying random erosion as corruption. Encoder inputs were provided as occupancy, with the target SDF estimated from the original meshes. For the corrupted dataset, the SDF was derived by converting the corrupted occupancies into a mesh and then estimating the SDF. The decoder follows the DeepSDF architecture with a dropout probability of $p=0.2$ where applicable. Quantitative results in Table 1 demonstrate that Dropsembles outperform other methods in terms of DSC, with EWC regularization providing further improvements in both DSC and ECE metrics across all methods. Although MC dropout achieves the lowest ECE score, it fails to preserve shape smoothness when used for uncertainty modeling rather than regularization, as highlighted by additional qualitative results in Appendix Figure 11.
Recently, non-ReLU activation functions have gained popularity in neural implicit architectures (Sitzmann et al., 2020; Tancik et al., 2020; Müller et al., 2022b; Mehta et al., 2021; Dupont et al., 2022). In Appendix Table 4, we demonstrate that Dropsembles can be effectively combined with periodic activation functions without any loss in performance.
@@
In this section, we adapt the data preparation procedure from (Turella et al., 2021) but replace the high-quality CT dataset with a synthetic dataset of anatomical shapes. As an anatomical shape prior, we utilize a rigged anatomical model from “TurboSquid”. To model patient-specific variability and variations in poses during MRI acquisition, we generated 94 rigged deformations. Point-cloud models from the atlas were converted to voxels at 256 voxel resolution in order to correspond to the resolution of the target MRI dataset. We further applied random elastic deformations to mimic patient-specific shape variability. We created a paired “sparse” - “dense” dataset by selecting a consistent set of 17–21 sagittal slices and additionally applied two iterations of connected erosions to simulate patient-specific automatic MR segmentations (Turella et al., 2021). A bicubic upsampling was employed on the “sparse” inputs to adapt them for the encoder. We trained ReconNet (Turella et al., 2021) on the entire training split of the atlas dataset to obtain an encoder, which we subsequently kept frozen. For the implicit decoder, we employed only one of the rigged samples to accurately quantify performance on synthetic data and demonstrate the method’s robustness to medium and strong distribution shifts. For real-world applications, we recommend using the entire rigged transformations dataset.
For the occupancy network, we employed the MLP architecture described by (Amiranashvili et al., 2022), which features eight linear layers, each of 128 dimensions and dropout $p=0.2$. The initial training of the occupancy network was conducted on an anatomical atlas. This intensive training process required 48 hours on a single NVIDIA RTX A6000 GPU, equipped with 48 GB of memory. This pretrained network was then used to fine-tune both Dropsembles and MC dropout models. Additionally, we trained four distinct instances of Ensembles without dropout, with each instance requiring 48 hours of training. A key distinction between Dropsembles and Ensembles is that Dropsembles leverage a single pretrained model from dataset A, significantly reducing computational demand by a factor equal to the number of thinned networks. This results in a substantial difference in total training time, as the pre-training on dataset A is much more time-consuming than the fine-tuning on dataset B (see Table 2 and Appendix E). Additionally, this training time on dataset A remains constant for Dropsembles as the number of thinned networks increases, while it grows linearly for Ensembles.
@@
As the target dataset B, we use a publicly available dataset of MR+CT images from 20 subjects (Cai et al., 2015). We employed the same segmentation network as (Turella et al., 2021) to obtain segmentations of 5 vertebrae, 5 discs, and the spinal canal (Pang et al., 2020). These automatically generated segmentations of high-quality MRI samples were used as the ground truth for the sparse 3D reconstruction task. To create sparse inputs, we removed the same set of sagittal slices as in the atlas dataset. We randomly selected 3 subjects for the consequent fine-tuning and testing.
The MR dataset described above lacks ground truth segmentations, which complicates numerical evaluation. The goal of the model is to impute missing or misclassified parts using anatomical atlas
@@
priors, which, while enhancing reconstruction, are typically marked as incorrect in standard segmentation metrics. To facilitate a more accurate assessment of reconstruction metrics, we introduce an intermediate benchmark. For this purpose, we utilize three random rigged deformations from the atlas that were not exposed during training and apply a doubled level of erosions. This strong augmentation provides a challenging input for dataset B and allows to model a significant distribution shift aimed at rigorously testing the model under adverse conditions.
Numerical evaluation in Table 2 demonstrates consistent improvement of our method upon MC-dropout and comparable performance to Ensembles. In order to investigate the performance of the proposed method on MR dataset, we perform a detailed qualitative analysis in Figure 4. A consistent and distinct pattern not captured by the metrics alone, stands out from the qualitative assessment: MC-dropout tends to produce noisy reconstructions, as illustrated in Figure 4c-d, and generally yields predictions that lack coherent, continuous shapes. This behaviour is partially captured in metric “DSC avg” in Table 2 - dice score evaluated on individual samples of the network instances. This observation is notable because, although models trained with conventional dropout generate reasonable predictions, their use in uncertainty estimation undermines the fundamental objective of 3D modeling. Samples drawn from the dropout distribution do not yield plausible shapes, as evidenced in Figure 4d. In contrast, our model, which draws from deep ensembles, does not exhibit these limitations. Furthermore, increasing the number of samples in MC-dropout does not resolve this issue but rather leads to higher computational demands, as documented in Table 2.
EWC demonstrates only minor improvements in terms of Dice score, while the difference in Hausdorff distance is more noticeable. We believe this is because Hausdorff distance better captures small discrepancies that the Dice score overlooks. Dice is less sensitive to small differences in reconstruction, as the majority of structures are well reconstructed by both methods. Qualitative inspection in Figure 5 confirms these subtle variations, which are crucial in medical applications where even small differences can significantly impact treatment planning.
Figure 5: Example in the performance difference between Dropsembles w/ and w/o EWC. Grey segmentations are produced by Dropsembles w/ or w/o EWC regularization overlayed with colorful ground-truth segmentations. EWC better captures subtle details in modeling thin structures, such as vertebra processes.
Table 2: Comparison of methods fine-tuned on dataset B of lumbar spine experiment. Evaluations are performed on the corrupted atlas. Metrics are reported for dataset B. Baseline is a MC-dropout model trained on dataset A (w/o fine-tuning).
Method DSC [%] $\uparrow$ DSC avg [%] $\uparrow$ HD $\downarrow$ ECE [%] $\downarrow$ Train-A / Tune-B [h]
Baseline $65.0\pm 2.0$ $63.9\pm 1.8$ $17.5\pm 1.2$ $25.2\pm 0.4$ 48 / -
MCdropout $84.8\pm 2.8$ $83.0\pm 2.0$ $18.15\pm 6.46$ $\underline{3.9}\pm 2.4$ 48 / 3.5
Dropsembles $\underline{86.8}\pm 2.4$ $\underline{86.3}\pm 2.3$ $11.4\pm 1.8$ $4.6\pm 1.7$ 48 / 10
Dropsembles + EWC $\textbf{86.9}\pm 2.4$ $\textbf{86.4}\pm 2.3$ $\underline{10.9}\pm 2.1$ $4.5\pm 1.7$ 48 / 11
Ensembles $86.5\pm 3.1$ $86.0\pm 2.9$ $\textbf{10.5}\pm 2.8$ $\textbf{3.9}\pm 2.2$ 192 / 10